OpenStudy (anonymous):

Please help me on this calculus In a model of epidemics, the number of infected individuals in a population at a time is a solution of the logistic differential equation dy/dt=0.6y-0.0002y^2 , where y is the number of infected individual in the community and t is the time in day. 1. Describe the population for this situation. 2. Assume that 10 people were infected at the initial time t=0 . Find the solution for the differential equation. 3. How many day will it take for half of the population to be infected?

OpenStudy (anonymous):

So we have: \[ \large \frac{dy}{dt} = 0.6y-0.0002y^2 \]

OpenStudy (anonymous):

It's seperable so:\[ \large \begin{array}{rcl} \frac{dy}{dt} &=& 0.6y-0.0002y^2 \\ \frac{1}{ 0.6y-0.0002y^2}\frac{dy}{dt} &=& 1 \\ \int \frac{1}{ 0.6y-0.0002y^2}\frac{dy}{dt} dt &=& \int 1 dt \\ \end{array} \]Remember that the differential is defined such that: \(dy = \frac{dy}{dt}dt\):\[ \int \frac{1}{ 0.6y-0.0002y^2} dy = \int 1 dt \]Now we just need to set up our limit of integration given initial conditions.

OpenStudy (anonymous):

Actually, scratch that, we'll just take the anti derivative of both and solve for the constants of integration later on.

OpenStudy (anonymous):

We're gonna want to use partial fractions, I think: \[ \begin{array}{rcl} \int \frac{1}{ 0.6y-0.0002y^2} dy &=& \int 1 dt \\ \int \frac{1}{ y(0.6-0.0002y)} dy &=& t+C_1 \end{array} \]

OpenStudy (anonymous):

We need to solve for \(A\) and \(B\):\[ \begin{array}{rcl} \frac{1}{ y(0.6-0.0002y)} &=& \frac{A}{y} + \frac{B}{0.6-0.0002y} \\ \frac{1}{ y(0.6-0.0002y)} &=& \frac{A(0.6-0.0002y)}{y(0.6-0.0002y)} + \frac{B(y)}{y(0.6-0.0002y)} \\ 1 &=& A(0.6-0.0002y) + B(y) \\ 1 &=& 0.6A-0.0002Ay + By \\ (1) + (0)y &=& 0.6A+(B-0.0002A)y \\ \end{array} \]This gives us a system of equations: \[ \begin{array}{rclrl} 1 &=& &0.6&A \\ 0 &=& B&-0.0002&A \end{array} \]

OpenStudy (anonymous):

@algradilla Think you can do the rest?

OpenStudy (anonymous):

I'm still confused on how I'm supposed to use the info that 10 people are infected when t=0